1. Express the following signals in the formy, plot the locations of Andy in the complex plane and indicate each signals sta(a)e-0.5, (b) 0.5k, (c) e jak, (c) ejak, (d) e-(+ja)k,

त्र -२ :(i
Answer #1

→ a) Given each signal to represent as F[k] = e = y* (t=e^] e-0.5% => 1 = -0.5 - (805)* »> t = -0.52 0.006s d-plane ↑ Im (a)As N =1 as k progresses the signal magnitude remains constant, but alternates. d) . estk x = in (eink » f = eit = -1 y-plane